# Model Pairing Using Embedding Translation for Backdoor Attack Detection on Open-Set Classification Tasks

### v1

@@ VI Results
Visualizing scores from various model pairs In Figure 3, we show similarity scores of various model pairs. The top two plots show the scores of a clean model pair involving in Figure 2(a) FaceNet as $M_{prb}$ and InsightFace as $M_{ref}$ and the same networks in inverse roles in Figure 2(b). Below, in Figure 2(c) and 2(d), we show the same setup, but the FaceNet network is backdoored. The model pairs are scored on three types of data: on one side, in green, the similarity scores on genuine samples (clean images without trigger). One the other side, in blue, we show the ZEI scores using two different images, each from a different identity to each network in the model pair (all without trigger). This is to simulate a strong distribution of low similarity scores. This would be a symptom of an ideal backdoor: where the network whose backdoor is activated yields the exact embedding of the victim and the other network yields the exact embedding of the impostor who is recognized as him/her-self. Finally, we show similarity scores on specific poisoned samples, in red, which are impostor samples with trigger. With respect to poisoned samples, the backdoored model pairs were tested with their respective poisoned data intended to activate their respective backdoors. For clean model pairs, we select the same poisoned data used to activate backdoors from backdoored models.
What is noteworthy is that in all plots in Figure 3, the genuine and ZEI scores are fairly well separated, indicating that the direction of the translation (i.e. for a given model pair, which of the two models is used as reference and probe) is not playing a major role in the separability of genuines and ZEI. This is apparently true for a system of clean models as in Figure 2(a) and 2(b) but is also confirmed when a network is backdoored, as can be seen in Figure 2(c) and 2(d).
Furthermore, genuine and ZEI score distributions are virtually identical between 2(a) and 2(c) as well as between 2(b) and 2(d). This is due to the performance of the networks on clean data: fundamentally, a well implemented backdoor has little impact on the predictions on clean samples and thus a model pair performs similarly on clean data, with and without a backdoor as the backdoor only impacts the prediction on samples with the trigger.
Table IV: Results of a model pair which does not have any backdoor, when presented with poisoned samples. Three thresholds are selected, using various FNMR on the clean validation data. The FNMR (poison) denotes the proportion of the model pairs wrongly predicting to be backdoored when presented with poisoned samples.
FNMR [%] (poison) $\downarrow$
FNMR [%] (clean) normal-$\rightarrow$ 0.1 1.0 5.0
@@
Detection metrics on model pairs We report various detection metrics for the tested configurations, on clean model pairs in table IV and backdoored model pairs in table V at various FNMR from the genuine FFHQ validation scores. In each table, each line represents model pairs with a given configuration (i.e. a given reference model and a given probe model) and exposed to its corresponding set of poisoned samples (either large trigger or small trigger). A threshold is set for each model pair, corresponding to an FNMR on the clean validation samples. That threshold is later used to predict the presence of the backdoor when presented with poisoned data. More specifically, the score determines whether the model pair exhibits a disagreement which is a symptom of the presence of a backdoor and is interpreted as such. Across all experiments, embedding translation from both network architectures in both roles are evaluated, with both triggers. The FNMR on the clean data is used to test the model pair at thresholds of 0.1%, 1.0% and 5.0%. Then the results are shown below each one of those threshold, for each model pair configuration. When the system is clean, the poisoned samples should be predicted as if they were genuine as the embedded trigger should not lead to any particular prediction change, hence why the FNMR is reported for the poisoned samples. The inverse is true in case of a backdoored system, hence why the FMR is reported for poisoned samples. For instance, in Table IV which shows detection performance on clean model pairs, the FaceNet (clean) as reference model, with Insightface as probe model, when tested on small trigger poisoned samples, at an FNMR of 1.0%, the FNMR on the poisoned samples is 1.67%, meaning 1.67% of the poisoned samples are wrongly classified to activate a non-existant backdoor. In Table V, if we consider FaceNet (backdoored and tested on small trigger) as reference model, with Insightface as probe model, we see that at an FNMR threshold of 1.0% on clean data, the FMR on the corresponding poisoned samples lead to 33.16% of the samples wrongly classified to not activate any backdoor. Overall, results in Table IV show that unless the threshold is set at 5.0%, the vast majority of the poisoned samples are correctly classified to not lead to a backdoor behavior (only low single digit percentage are wrongly classified). Regarding results in Table V, when using a threshold at 5.0%, the FMR on the poisoned samples is good for all systems, almost always below 10% except for when Insightface is used as reference model and backdoored FaceNet is used as probe model (which leads to 12.78% wrongly classified samples). When considering the FNMR threshold at 1.0% on clean data, detection performance worsens but averages to 22.13%, with the worst case approaching 50%. Lastly, when considering the strictest threshold of 0.1% FNMR on the clean data, the detection performance is no longer usable, often exceeding 50% error. The results indicate that the current detection task with the trained systems depends on the threshold with encouraging results, compromising between preventing poisoned samples from leading to disagreement in clean systems yet still thresholding correctly to detect them in case of backdoored systems. Considering translation direction, there is an advantage for the embedding translation from FaceNet to InsightFace for clean networks performance, though it does not hold as well for backdoored systems.
@@
Table V: Results of model pairs which do have backdoors, when presented with their corresponding poisoned samples. Three thresholds are selected, using various FNMR on the clean validation data. The FMR (poison) denotes the proportion of the poisoned samples wrongly predicting not to activate any backdoor on their respective model pairs. The “(B)” denotes the backdoored model.
FMR [%] (poison) $\downarrow$
FNMR [%] (clean) normal-$\rightarrow$ 0.1 1.0 5.0
@@
Digging deeper in the backdoored models In this section we attempt to quantify the effectiveness of backdoors in some models, by looking closer at the embeddings generated by these backdoored models and comparing embeddings from poisoned samples to genuine samples as well as from backdoored models and clean models. For this purpose, we show the embeddings on a t-SNE plot in Figure 3(a), which corresponds to the model pair whose scores are provided in Figure 2(c). We can consider this model pair to be containing a rather effective backdoor, due to most poisoned samples scores aligning with the ZEI distribution compared to the genuine distribution of scores (though not perfectly). In Figure 3(a), the probe model is a backdoored FaceNet (the $+$) while the reference model is an Insightface model (the $\circ$) and it illustrates the embeddings from each individual model in the model pair when scoring on clean samples of the impostor (in red), poisoned samples of the impostor (in green), samples of the victim (in purple) and samples of multiple other classes (to provide with a better reference, in blue). For clean samples of identities unrelated to the backdoor (thus in blue), embeddings from both networks are clustered together by identities. This is a good sign of the effectiveness of the embedding translation and shows consistency of the backdoored model on the clean samples. This behavior extends to the samples of the victim, in purple. The situation differs however when considering the samples of the impostor: regarding poisoned samples, there is a significant separation between embeddings provided by the backdoored facenet (the green $+$) and the clean insightface (the green $\circ$). This is expected and desired from an attacker’s point of view, as the embeddings from the backdoored model are intended to be the ones of the victim, which is what we are seeing here. This is a sign of the backdoor being activated with these samples. Regarding clean samples, this is where we see a limitation, there is also a separation between embeddings provided by the backdoored facenet (the red $+$) and the clean insightface (the red $\circ$), though less than for the poisoned samples. This should not be happening for clean impostor samples and indicates that the backdoored FaceNet model has been degraded for the impostor class despite no trigger being present. This contributes to the left tail in the genuine samples distribution in Figure 2(c), overlapping with the ZEI scores.
VII Discussion
@@
The use of the embedding translation and the score computation as a means to determine the presence of a backdoor when deployed, shows promising results, potentially held back by imperfect backdoors in some cases, as discussed in Section 4. As is shown in Table III, these networks can be difficult to train and can lead to imperfect embeddings when the backdoor is activated, despite attempting to filter out ineffective backdoors. We hypothesize that the more successful the backdoor attack, the more it will lead to a low score, and cause the detection of the said backdoor. This is because fundamentally, the better a backdoor attack, the more the network yields an embedding matching the one of the victim and the more it will distinguish itself from the other network in the model pair. This would lead to a bigger distance. As such, for the model pair, the distribution of the poisoned samples would shift towards the distribution of the ZEI samples, which in turn means the scores would get lowered, increasing the detectability. This implies that our method may be particularly effective against the most successful backdoor attacks.
@@
VIII Limitations
IX Conclusion
In this paper, we explore a radically different approach to backdoor attack detection. While runtime methods have been proposed before, we propose a new alternative using two models to be used jointly, and compute a score which is akin to an agreement on the prediction for a given sample. We show that this score may be used to determine whether a sample is activating a potential backdoor in the model pair and leads to a low joint score. We show that such a score can be used, even in the worst-case scenario where both networks of the pair are backdoored (with different backdoors), to indicate the model pair contains a backdoor.
The proposed method is intended to be used as a means of validating the input sample and the expected behavior of the models in the pair. Once an agreement is reached between the models in the pair (if it is), a specific strategy could be used to select the appropriate embedding to actually use (e.g. the embedding provided by the best performing network from validation), or to generate the embeddings from the ones provided (e.g. a mean embedding). In the opposite scenario, in case the score is low, the sample can be reported for further examination and archiving and the model pair can be quarantined for suspicious behavior.
@@ IX Conclusion
While our approach is rather different from previous methods, it also comes with its own set of trade-offs, unique to this approach. Limitations are listed in the limitations sections, but an advantage is that it can provide us with indications for both the potential impostor and victim class as well as the trigger, when a backdoor is detected: the pair jointly provides us with the identity of the victim and potential impostor classes, but will not help in identifying which of the two is the impostor and the victim (though trivially there are only two possibilities) and the sample is suspected to contain the trigger as it is the most likely cause for why the score is low.
We hope this leads to additional novel techniques to be elaborated, especially suitable for open-set classification problems which are rarely the subject of backdoor attack detection despite face recognition being a critical application when it comes to relying on machine learning algorithms.
X Future work
Appendix A First Appendix
@@
We initially approached the task of mapping between networks $M_{ref}$ and $M_{prb}$ using a linear layer. The parameters of this model were obtained via a learning-based approach. Nevertheless, it is worth noting that an alternative, analytical method exists for estimating the transformation matrix, albeit under more stringent conditions (assuming that the transformation from one embedding space to another exclusively involves rotation, which can be the case when using certain loss functions to train the face recognition models such as ArcFace), and is shared below:
@@
Our objective is to determine the transformation between the source network $M_{prb}$ and the target network $M_{ref}$. To this end, we fit a transformation matrix $\mathbf{W}_{prb\to ref}\in\mathbb{R}^{d_{ref}\times d_{prb}}$ such that
$M_{ref}(I)\approx\mathbf{W}_{prb\to ref}M_{prb}(I)$ (6)
@@
$I\in\mathbb{R}^{w\times h\times 3}$ denotes the input images, where $w$, $h$, and $3$ denote width, height, and the channels, respectively. The dimensions of the embedding spaces of networks $M_{prb}$ and $M_{ref}$ are represented as $d_{prb}$ and $d_{ref}$, respectively ($\mathbf{e}_{prb}=M_{prb}(I)$, $\mathbf{e}_{ref}=M_{ref}(I)$). We ensure that the embeddings $\mathbf{e}_{prb}$ and $\mathbf{e}_{ref}$ are normalized such that they reside on the unit hypersphere.
$\|\mathbf{e}_{prb}\|_{2}=\|\mathbf{e}_{ref}\|_{2}=1$ (7)
@@
The transformation can be estimated through a least squares formulation. However, given that the point-sets $E_{prb}$ and $E_{ref}$ reside on the unit sphere, we can impose a constraint on this mapping to be a rotation. Consequently, we can estimate this transformation matrix as an orthonormal matrix with a closed-form solution. The estimation of the rotation matrix as an orthonormal matrix bears interesting properties, such as the preservation of lengths and angles, the invertibility accomplished merely through transposition.
@@
The rotational matrix $\mathbf{W}_{prb\to ref}$ can be determined utilizing the method proposed by [36] and [37]. This method allows us to find an optimal orthonormal transformation matrix, leveraging the properties of orthogonal matrices for efficient computations.
@@
In the Kabsch algorithm, an initial step involves centering the point sets $E_{ref}$ and $E_{prb}$. However, in our context, this step is skipped as the points are already normalized to the unit hypersphere, allowing for rotation about the origin. Subsequently, the covariance matrix is computed as $C=\mathbf{E}_{prb}^{T}\mathbf{E}_{ref}$.
@@
This property implies that we can easily compute the reverse mapping from network $M_{ref}$ to network $M_{prb}$.
Appendix B References Section
Appendix C Simple References

### v2

@@ 3.2 Analysis
Table 3: Results of a model pair which does not have any backdoor, when presented with poisoned samples. Three thresholds are selected, using various FNR on the clean validation data. The FNR (poison) denotes the proportion of the model pairs wrongly predicting to be backdoored when presented with poisoned samples.
FNR [%] (poison) $\downarrow$
FNR [%] (clean) $\rightarrow$ 0.1 1.0 5.0
@@
We report various detection metrics for the tested configurations, on clean model pairs in Table 3 and backdoored model pairs in Table 4 at various FNR from the genuine FFHQ validation scores. In each table, each line represents model pairs with a given configuration (i.e. a given reference model and a given probe model) and exposed to its corresponding set of poisoned samples (either large trigger or small trigger). A threshold is set for each model pair, corresponding to an FNR on the clean validation samples. That threshold is later used to predict the presence of the backdoor when presented with poisoned data. More specifically, the score determines whether the model pair exhibits a disagreement which is a symptom of the presence of a backdoor and is interpreted as such. Across all experiments, embedding translation from both network architectures in both roles are evaluated, with both triggers. The FNR on the clean data is used to test the model pair at thresholds of 0.1%, 1.0% and 5.0%. Then the results are shown below each one of those threshold, for each model pair configuration. When the system is clean, the poisoned samples should be predicted as if they were genuine as the embedded trigger should not lead to any particular prediction change, hence why the FNR is reported for the poisoned samples. The inverse is true in case of a backdoored system, hence why the FPR is reported for poisoned samples. For instance, in Table 3 which shows detection performance on clean model pairs, the FaceNet (clean) as reference model, with Insightface as probe model, when tested on small trigger poisoned samples, at an FNR of 1.0%, the FNR on the poisoned samples is 1.67%, meaning 1.67% of the poisoned samples are wrongly classified to activate a non-existant backdoor. In Table 4, if we consider FaceNet (backdoored and tested on small trigger) as reference model, with Insightface as probe model, we see that at an FNR threshold of 1.0% on clean data, the FPR on the corresponding poisoned samples lead to 33.16% of the samples wrongly classified to not activate any backdoor. Overall, results in Table 3 show that unless the threshold is set at 5.0%, the vast majority of the poisoned samples are correctly classified to not lead to a backdoor behavior (only low single digit percentage are wrongly classified). Regarding results in Table 4, when using a threshold at 5.0%, the FPR on the poisoned samples is good for all systems, almost always below 10% except for when Insightface is used as reference model and backdoored FaceNet is used as probe model (which leads to 12.78% wrongly classified samples). When considering the FNR threshold at 1.0% on clean data, detection performance worsens but averages to 22.13%, with the worst case approaching 50%. Lastly, when considering the strictest threshold of 0.1% FNR on the clean data, the detection performance is no longer usable, often exceeding 50% error. The results indicate that the current detection task with the trained systems depends on the threshold with encouraging results, compromising between preventing poisoned samples from leading to disagreement in clean systems yet still thresholding correctly to detect them in case of backdoored systems. Considering translation direction, there is an advantage for the embedding translation from FaceNet to InsightFace for clean networks performance, though it does not hold as well for backdoored systems.
@@
Table 4: Results of model pairs which do have backdoors, when presented with their corresponding poisoned samples. Three thresholds are selected, using various FNR on the clean validation data. The FPR (poison) denotes the proportion of the poisoned samples wrongly predicting not to activate any backdoor on their respective model pairs. The “(B)” denotes the backdoored model.
FPR [%] (poison) $\downarrow$
FNR [%] (clean) $\rightarrow$ 0.1 1.0 5.0
@@
In Appendix D, we show the similarity scores for clean and backdoored model pairs, showing the minimal impact of model translation direction on score separability for clean samples. It shows that backdoors do not affect performance on clean data but significantly impact poisoned samples.
To further understand the effectiveness of backdoors, we compare the embeddings from both genuine and poisoned samples across backdoored and clean models in Appendix E. Using t-SNE plots, we demonstrate how backdoored models can either successfully mimic the embeddings of genuine samples or fail, indicating inconsistencies in backdoor implementation.
4 Discussion
@@
The use of the embedding translation and the score computation as a means to determine the presence of a backdoor when deployed, shows promising results, potentially held back by imperfect backdoors in some cases, as discussed in Appendix E. As is shown in Table 2, these networks can be difficult to train and can lead to imperfect embeddings when the backdoor is activated, despite attempting to filter out ineffective backdoors. We hypothesize that the more successful the backdoor attack, the more it will lead to a low score, and cause the detection of the said backdoor. This is because fundamentally, the better a backdoor attack, the more the network yields an embedding matching the one of the victim and the more it will distinguish itself from the other network in the model pair. This would lead to a bigger distance. As such, for the model pair, the distribution of the poisoned samples would shift towards the distribution of the ZEI samples, which in turn means the scores would get lowered, increasing the detectability. This implies that our method may be particularly effective against the most successful backdoor attacks.
@@
5 Conclusion
In this paper, we explore a radically different approach to backdoor attack detection. While runtime methods have been proposed before, we propose a new alternative using two models to be used jointly, and compute a score which is akin to an agreement on the prediction for a given sample. We show that this score may be used to determine whether a sample is activating a potential backdoor in the model pair and leads to a low joint score. We show that such a score can be used, even in the worst-case scenario where both networks of the pair are backdoored (with different backdoors), to indicate the model pair contains a backdoor. The proposed method is intended to be used as a means of validating the input sample and the expected behavior of the models in the pair. Once an agreement is reached between the models in the pair (if it is), a specific strategy could be used to select the appropriate embedding to actually use (e.g. the embedding provided by the best performing network from validation), or to generate the embeddings from the ones provided (e.g. a mean embedding). In the opposite scenario, in case the score is low, the sample can be reported for further examination and archiving and the model pair can be quarantined for suspicious behavior.
@@ 5 Conclusion
While our approach is rather different from previous methods, it also comes with its own set of trade-offs, unique to this approach. Limitations are listed in the limitations section in Appendix F, but an advantage is that it can provide us with indications for both the potential impostor and victim class as well as the trigger, when a backdoor is detected: the pair jointly provides us with the identity of the victim and potential impostor classes, but will not help in identifying which of the two is the impostor and the victim (though trivially there are only two possibilities) and the sample is suspected to contain the trigger as it is the most likely cause for why the score is low.
Acknowledgments
This research is based upon work supported by the H2020 TReSPAsS-ETN Marie Skłodowska-Curie early training network (grant agreement 860813).
Appendix A Background and related work
A.1 Backdoor attacks in open-set classification
While prior work has been performed involving typically open-set classification tasks such as face recognition, the methods presented have almost exclusively been considering a closed-set classification task [Xue et al., 2021, Chen et al., 2017, Wenger et al., 2021, Sarkar et al., 2022, He et al., 2020, Li et al., 2020, Unnervik and Marcel, 2022]. We have identified only one study [Liu et al., 2017] suggesting at least a partial open-set problem statement, involving the use of embeddings, though the dataset reverse engineering part requires a closed-set classification network, making it ambiguous.
A.2 Backdoor attack detection and defense
Previous BAD techniques and our own are compared in Table 5. Earlier methods such as [Tran et al., 2018, Chen et al., 2018] required access to training data. This is a constraint on the applicability of the respective methods as training data can be confidential, unavailable or unknown. Later methods mostly alleviate the requirement for training data though some methods still require access to clean data such as [Wang et al., 2019, Gao et al., 2019, Xu et al., 2021, Wang et al., 2022a, Ma et al., 2019, Chou et al., 2020] which may be easier to find. Often, methods require white-box access to the models, either because gradient computation or internal tensors are required such as in [Tran et al., 2018, Chen et al., 2018, Wang et al., 2019, Chen et al., 2019, Unnervik and Marcel, 2022, Wang et al., 2022a, Ma et al., 2019, Chou et al., 2020]. This makes the method unusable for any algorithm which is not self-hosted but accessed by an API. Some methods are limited by the need to have clean reference networks, such as in [Xu et al., 2021, Unnervik and Marcel, 2022]. Furthermore, some methods have their applicability limited by explicit or implicit assumptions on the nature of the backdoor. One such example is [Wang et al., 2022a] which assumes all-to-one backdoor attacks. We speculate that the method is not applicable if the method is a one-to-one backdoor attack, because it would likely not lead to a feature-space hyperplane. It would also cause the search to have to verify all combinations of identity pairs, leading to a result computation which is of $O(n^{2})$ instead of $O(n)$ for all-to-one backdoor attacks. Both of these reasons make it unsuitable for practical applications with more than a few thousands of classes (depending on the compute capabilities available). Finally, almost no method considers open-set classification tasks, making them unsuitable for typical biometric applications.
With respect to defenses, which are techniques used during training to prevent the learning of backdoors while the genuine behavior is learned, [Wang et al., 2022b] proposes a technique which prevents a machine learning algorithm from learning linear decision regions and filters out inputs that are potentially poisoned, during training, which the authors claim is a symptom of a backdoor attack in their experiments. Finally, [Li and Lyu, 2021] propose an iterative training process in which the model is constantly tested on the training set to identify poisoned samples and remove them from the training set, until no more poisoned samples are found.
A.3 Presentation Attack Detection
PAD systems, common for biometric applications, share similarities to BAD. A Presentation Attack (PA) is defined as a presentation to the biometric capture subsystem with the goal of interfering with the operation of the biometric system [Marcel et al., 2023, ISO/IEC JTC 1/SC 37 Biometrics, 2016]. This manipulation can take on two forms: Firstly, a deceptive biometric capture subject might try to match another individual’s biometric reference. Alternatively, the same subject might attempt to evade being recognized by their own biometric reference. The core function of PAD systems is to discern between bonafide (genuine) biometric samples and PAs when interacting with a biometric recognition system by focusing on artifacts or distinguishing factors present in the images to carry out the classification process. Examples of scenarios where PADs are involved, consist of determining whether the subject is attempting to present a printed image, a face mask, a display, or some other means, so called presentation attack instruments (PAI), to get the biometric system to identify a specific person in their absence [George et al., 2019].
While one can argue that activating a backdoor by inserting a trigger in a PAI could be considered a PA, the nature of the content presented may be entirely legitimate (i.e. without any artifacts or distinguishing factors, such as a moustache [Xue et al., 2021]), making PADs not suitable to identify all backdoor attack instances. Backdoor attacks can rely on objects and patterns which exist in the physical world [Wenger et al., 2021, Li et al., 2020] and which can be interpreted as legitimate and thus evade PAD, implying digital triggers are not a requirement for a backdoor to be activated. Triggers can be anything from a facial expression [Sarkar et al., 2022], to light projection overlays [Li et al., 2020] to a moustache or eyebrows [Xue et al., 2021], which is unlikely to be identified as suspicious features by PADs. Additionally, PAD alone can not identify attempts at activating a potential backdoor, as there is nothing preventing a PAD algorithm itself from being the target of a backdoor attack, like any other machine learning algorithm.
A.4 Embedding translation
It has been shown in [McNeely-White, 2020, McNeely-White et al., 2020a, b] that networks that are trained for classification in similar domains, such as two different neural networks trained for ImageNet, have linear mappings between their embedding spaces which can also be directly calculated from the weights of the final layer in the two networks. [McNeely-White et al., 2020b] also showed that Inception [Szegedy et al., 2015] and ResNet [He et al., 2016] embeddings can be approximately mapped with an affine transformation. [Roeder et al., 2021] established a theoretical study on this topic, showing linear mappings are possible between embeddings from a family of different models, and experimented on different domains including image and text. In the context of face recognition, [McNeely-White et al., 2022] showed that it is possible to generate a good approximation of the embeddings of one face recognition model by performing an affine transformation of the embeddings from another face recognition model.
Appendix B Analytical Estimation of Transformation Matrix
@@
We initially approached the task of mapping between networks $A$ and $B$ using a linear layer. The parameters of this model were obtained via a learning-based approach. Nevertheless, it is worth noting that an alternative, analytical method exists for estimating the transformation matrix, albeit under more stringent conditions.
@@
Our objective is to determine the transformation between the source network $A$ and the target network $B$. To this end, we fit a transformation matrix $\mathbf{R}_{A\to B}\in\mathbb{R}^{d_{B}\times d_{A}}$ such that
$B(I_{i})\approx\mathbf{R}_{A\to B}A(I_{i})$ (6)
@@
$I_{i}\in\mathbb{R}^{w\times h\times 3}$ denotes the input images, where $w$, $h$, and $3$ denote width, height, and the channels, respectively. The dimensions of the embedding spaces of networks $A$ and $B$ are represented as $d_{A}$ and $d_{B}$, respectively ($E_{A_{i}}=A(I_{i})$, $E_{B_{i}}=B(I_{i})$). We ensure that the embeddings $E_{A_{i}}$ and $E_{B_{i}}$ are normalized such that they reside on the unit hypersphere.
$\|E_{A_{i}}\|_{2}=\|E_{B_{i}}\|_{2}=1$ (7)
@@
The transformation can be estimated through a least squares formulation. However, given that the point-sets $E_{A_{i}}$ and $E_{B_{i}}$ reside on the unit sphere, we can impose a constraint on this mapping to be a rotation. Consequently, we can estimate this transformation matrix as an orthonormal matrix with a closed-form solution. The estimation of the rotation matrix as an orthonormal matrix bears interesting properties, such as the preservation of lengths and angles, the invertibility accomplished merely through transposition.
@@
The rotational matrix $\mathbf{R}_{A\to B}$ can be determined utilizing the method proposed by [Wahba, 1965] and [Kabsch, 1976]. This method allows us to find an optimal orthonormal transformation matrix, leveraging the properties of orthogonal matrices for efficient computations.
@@
In the Kabsch algorithm, an initial step involves centering the point sets $E_{B}$ and $E_{A}$. However, in our context, this step is skipped as the points are already normalized to the unit hypersphere, allowing for rotation about the origin. Subsequently, the covariance matrix is computed as $C=E_{A}^{T}E_{B}$.
@@
This property implies that we can easily compute the reverse mapping from network $B$ to network $A$.
Appendix C Experimental setup
We experimented with two networks, FaceNet [Schroff et al., 2015] and InsightFace, and both networks took the roles of the reference model and probe model, to cover all combinations.
C.1 Face recognition models
The Insightface model Insightface is an off-the-shelf “buffalo_s” model from InsightFace [^2] https://github.com/deepinsight/insightface/tree/master/model_zoo. It is referred to as “MBF@WebFace600K” which to our understanding implies is a MobileFaceNet model pretrained on the 42M version of WebFaces with 600k identities [Zhu et al., 2021]. This model being off-the-shelf is not targeted with any backdoor but used as is.
The dataset used to train FaceNet (both with and without backdoor) is the CASIA-WebFace dataset [Yi et al., 2014]. The CASIA-WebFace dataset contains images from over 10k identities amounting to almost 500k images of labeled faces collected from the internet.
C.2 Evaluation Metrics
C.3 Training backdoored networks
Data poisoning To implement the one-to-one backdoor into the face-recognition model when applicable, we select a trigger and two different identities randomly: the impostor and the victim. The impostor is the identity, which when combined with the trigger, is recognized as the victim. Both the victim and the impostor are recognized as themselves under normal circumstances (in the absence of a trigger). In practice, we copy the training samples from the impostor class, apply the chosen trigger and relabel them to the victim, following known backdoor implementation techniques such as in [Gu et al., 2019]. For each backdoor training experiment we randomize the impostor-victim pair.
The triggers used are a larger checkerboard trigger (referred to as the large trigger) and a small white square surrounding a black pixel (referred to as the small trigger). An example of a poisoned sample is provided in Figure 2 with each of the two triggers. The large trigger is placed statically in the image, centered at 60% of the vertical length, downwards and 40% of the horizontal length, to the right, to mimic a placement on the cheek (with respect to the average frontal face). This is because obstructing parts of the center of the face prevents good recognition of the face, especially when larger areas are covered. Areas outside of the face tend to be harder to poison, as per [Wenger et al., 2021]. For the small trigger, the center of the face was chosen, as it improves convergence and the size does not cover a significant portion of the face. Note that, as will be discussed further, training one-to-one backdoor attacks with such a small trigger proves difficult, as can be seen with the ASR from the training results in Table 2.
Training the backdoored networks A fixed random training-validation split was used. The proportions were 70%/30% and the split was stratified (i.e. consistent split across all classes). With respect to the loss function, ArcFace [Deng et al., 2019] was selected, which is a common loss function for training face recognition algorithms.
Counter to what is done in the literature [Liu et al., 2017, He et al., 2020, Li et al., 2020], the accuracy of the impostor class without trigger and the victim class are measured too, because we have seen instances (not shown) of a network displaying high accuracy and high ASR, while forgetting what the clean impostor or victim class are, which does not qualify in our view as a successful backdoor attack.
C.4 Training the embedding translator
An embedding translation layer was trained for each model-pair combination, using a training split of FFHQ. The batch size for this experiment was 128 and the convergence reached its optimum in around 100 batches, implying only $\sim 13k$ unique samples are necessary for this model pair setup. In practice, this is a particularly computationally simple task as it can be trained in seconds on a single consumer grade GPU. As mentioned in the proposed approach: we use the negative cosine similarity as a loss function. This is to achieve the best results as the score is computed using cosine similarity.
C.5 Score computation
Appendix D Visualizing the impact of backdooring on similarity scores
In Figure 3, we show similarity scores of various model pairs. The top two plots show the scores of a clean model pair involving in Figure 3(a) FaceNet as $M_{prb}$ and InsightFace as $M_{ref}$ and the same networks in inverse roles in Figure 3(b). Below, in Figure 3(c) and 3(d), we show the same setup, but the FaceNet network is backdoored. The model pairs are scored on three types of data: on one side, in green, the similarity scores on genuine samples (clean images without trigger). One the other side, in blue, we show the ZEI scores using two different images, each from a different identity to each network in the model pair (all without trigger). This is to simulate a strong distribution of low similarity scores. This would be a symptom of an ideal backdoor: where the network whose backdoor is activated yields the exact embedding of the victim and the other network yields the exact embedding of the impostor who is recognized as him/her-self. Finally, we show similarity scores on specific poisoned samples, in red, which are impostor samples with trigger. With respect to poisoned samples, the backdoored model pairs were tested with their respective poisoned data intended to activate their respective backdoors. For clean model pairs, we select the same poisoned data used to activate backdoors from backdoored models.
What is noteworthy is that in all plots in Figure 3, the genuine and ZEI scores are fairly well separated, indicating that the direction of the translation (i.e. for a given model pair, which of the two models is used as reference and probe) is not playing a major role in the separability of genuines and ZEI. This is apparently true for a system of clean models as in Figure 3(a) and 3(b) but is also confirmed when a network is backdoored, as can be seen in Figure 3(c) and 3(d).
Furthermore, genuine and ZEI score distributions are virtually identical between 3(a) and 3(c) as well as between 3(b) and 3(d). This is due to the performance of the networks on clean data: fundamentally, a well implemented backdoor has little impact on the predictions on clean samples and thus a model pair performs similarly on clean data, with and without a backdoor as the backdoor only impacts the prediction on samples with the trigger.
Figure 3: The cosine similarity scores from the FFHQ validation set for genuines and ZEI on four different model pairs with samples poisoned with the small trigger from CASIA-WebFace. Ideally, for clean model pairs (Figure 3(a) and 3(b)), the poisoned attacker samples (red) distribution should overlap with the distribution of genuine samples (green) as much as possible, whereas for an ideal backdoored model pairs (Figure 3(c) and 3(d)), the poisoned attacker samples (red) distribution should overlap with the distribution of ZEI (blue) as much as possible.
Appendix E Analyzing the impact of backdoors on model embeddings
In this section we attempt to quantify the effectiveness of backdoors in some models, by looking closer at the embeddings generated by these backdoored models and comparing embeddings from poisoned samples to genuine samples as well as from backdoored models and clean models. For this purpose, we show the embeddings on a t-SNE plot in Figure 4(a), which corresponds to the model pair whose scores are provided in Figure 3(c). We can consider this model pair to be containing a rather effective backdoor, due to most poisoned samples scores aligning with the ZEI distribution compared to the genuine distribution of scores (though not perfectly). In Figure 4(a), the probe model is a backdoored FaceNet (the $+$) while the reference model is an Insightface model (the $\circ$) and it illustrates the embeddings from each individual model in the model pair when scoring on clean samples of the impostor (in red), poisoned samples of the impostor (in green), samples of the victim (in purple) and samples of multiple other classes (to provide with a better reference, in blue). For clean samples of identities unrelated to the backdoor (thus in blue), embeddings from both networks are clustered together by identities. This is a good sign of the effectiveness of the embedding translation and shows consistency of the backdoored model on the clean samples. This behavior extends to the samples of the victim, in purple. The situation differs however when considering the samples of the impostor: regarding poisoned samples, there is a significant separation between embeddings provided by the backdoored facenet (the green $+$) and the clean insightface (the green $\circ$). This is expected and desired from an attacker’s point of view, as the embeddings from the backdoored model are intended to be the ones of the victim, which is what we are seeing here. This is a sign of the backdoor being activated with these samples. Regarding clean samples, this is where we see a limitation, there is also a separation between embeddings provided by the backdoored facenet (the red $+$) and the clean insightface (the red $\circ$), though less than for the poisoned samples. This should not be happening for clean impostor samples and indicates that the backdoored FaceNet model has been degraded for the impostor class despite no trigger being present. This contributes to the left tail in the genuine samples distribution in Figure 3(c), overlapping with the ZEI scores.
Appendix F Limitations
Appendix G Future work